# Metric theory of inhomogeneous Diophantine approximations with a fixed matrix

### v2

@@ 7.2 Proof of Theorem 1
Let us consider a decreasing function $\psi$ and function $\rho$ defined by (26); notice that such $\psi$ and $\rho$ are connected by (29). Let
@@
$\sigma(t)=\frac{\varepsilon}{n\rho\left(\frac{m}{\varepsilon}\cdot t\right)}.$ (43)
@@
Consider the infinite increasing sequence of integer vectors ${\bf y}\in\mathbb{Z}^{n}$ for which (42) holds. We may assume that this sequence is a subsequence of the best approximation vectors for matrix $\Theta^{\top}$. To avoid cumbersome notation, in this proof we do not denote this subsequence as ${\bf y}_{\nu_{k}},k=1,2,3,$..., but simply as ${\bf y}_{\nu},\nu=1,2,3,$... and do not use double indices. Let $Y_{\nu}=|{\bf y}_{\nu}|$. Let
$t_{\nu}=\frac{\varepsilon}{m\psi(Y_{\nu})}.$ (44)
It is clear that $t_{\nu}\to\infty$ when $\nu\to\infty$. Therefore, for $\boldsymbol{\eta}\in\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\mathcal{A}$ and for large enough $\nu$ there exists such a ${\bf q}\in\mathbb{Z}^{m}$ that $|{\bf q}|\leq t_{\nu}$ and $\Theta{\bf q}-\boldsymbol{\eta}||\leq\sigma(t_{\nu})$.
Now by (34), one has
@@
Note that by (44) we have
@@
Now the last two equalities by (45) imply
@@
From (46) we see that
$\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\mathcal{A}\cap\left(B_{M}(0)+\mathfrak{a}\right)\subseteq\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}\Omega_{\nu}$ (47)
(recall that now $B_{M}(0)+\mathfrak{a}$ is a ${\rm dim}\,\mathcal{A}$-dimensional ball of radius $M$ in sup-norm in $\mathcal{A}$ centered at $\mathfrak{a}\in\mathcal{A}$), so it is enough to prove that for any fixed $M$ and $\mathfrak{a}$ the latter set has measure zero under given restrictions on $\varepsilon$. Further, by (47) we have
$\left(B_{M}(0)+\mathfrak{a}\right)\setminus\widehat{\bf D}_{n,m}^{\Theta}[\tilde{g}]\supset\left(B_{M}(0)+\mathfrak{a}\right)\setminus\left(\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}\Omega_{\nu}\right)=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{\nu=N}^{\infty}\Omega_{\nu}^{c},\,\,\,\text{where}\,\,\,\Omega_{\nu}^{c}=\left(B_{M}(0)+\mathfrak{a}\right)\setminus\Omega_{\nu}.$
Now, we should note that by the assumption of Theorem 1, $\mathcal{A}$ is non-$\Xi$-exceptional for some $\Xi<1$. This means that (24) is not satisfied an so $|{\rm pr}_{\mathcal{L}}\,{\bf y}_{\nu}|\rightarrow\infty$. By Lemma 4 we see that
$m({\Omega}_{\nu})\leq 4\varepsilon(1+\xi(M))\cdot m(B_{M}(0)),$ (48)
where $\lim\limits_{M\rightarrow\infty}\xi(M)=0$. As $0<\varepsilon<\frac{1}{4}$, by (48) we can take $M$ large enough to ensure the inequality $4\varepsilon(1+\xi(M))<1$. Now
$m(\tilde{\Omega}_{\nu}^{c})\geq\left(1-4\varepsilon(1+\xi(M))\right)m(B_{M}(0))>Cm(B_{M}(0))$
@@
$\sum\limits_{k=1}^{\infty}m(\tilde{\Omega}_{\nu}^{c})=\infty.$ (49)
As we have noticed in the beginning of our proof, we may take a subsequence of the sequence of all the best approximations. Now it is necessary to take even more rare subsequence
to satisfy the condition
@@
Under this condition we can apply Lemma 5. Together with Lemma J and (49) this gives
@@
$m\left(\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\left(B_{M}(0)+\mathfrak{a}\right)\right)\leq m\left(\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}{\Omega}_{\nu}\right)=m(B_{M}(0))-m(B_{M}(0))=0,$
@@
$\{r:\,\,1\leq r\leq W_{k}^{\prime},\,\,I_{s}({\bf q}^{l}(s))\cap I_{k}({\bf q}^{r}(k))\neq\emptyset\}\leq\Big{\lceil}\frac{2\kappa\psi_{s}\cdot M_{k}}{2\cdot M_{s}\cdot 2\kappa}+\frac{1}{2}\Big{\rceil}\leq(1+\varepsilon)W_{k}^{\prime}m\left(I_{s}({\bf q}^{l}(s))\right)$ (51)
@@
which implies that $\sum\limits_{k=1}^{\infty}m(E_{k})=\infty$. We also know by (51) that for any $s<k$,
@@
The next statement directly implies Theorem 3 via the identification (26).
@@
$\text{the inequality}\,\,\,\psi_{\Theta^{\top}}(t)>\psi(t)\,\,\,\text{holds for an unbounded set of $t\in\mathbb{R}_{+}$.}$ (52)
@@
First of all, let us notice that the condition (52) implies that $\Theta^{\top}$ is not trivially singular (otherwise, $\psi_{\Theta^{\top}}(t)$ would become zero for all large enough $t$), so the sequence of the best approximation vectors for $\Theta^{\top}$ is well defined. Take any $\{\psi_{k}\}$ as in Lemma 8. Let $\{\nu_{k}\}$ be a sequence of natural numbers, such that $\psi_{\Theta^{\top}}(|{\bf y}_{\nu_{k}}|)\geq\psi(Y_{\nu_{k}+1})$, and rare enough to guarantee $\psi_{k}\frac{Y_{\nu_{k}+1}}{Y_{\nu_{k}}}\rightarrow\infty$.
@@
8 Proof of Theorem 7 and Theorem 4 C
We want to deduce Theorem 7 and Theorem 4 C from Theorem 6. We will only show the deduction of Theorem 7; the proof of Theorem 4 C is purely technical and follows the same lines.
First, note that $\lambda(T):=\left(\frac{u(T)}{g(T)}\right)^{\frac{m\gamma}{m+\gamma n}}$. This form implies that the condition (13) holds.
Properties (8) and (9) imply that the condition (14) holds: suppose $\delta>1$ and $T\in\mathbb{R}$, then
@@ 8 Proof of Theorem 7 and Theorem 4 C
Thus, the function $\tilde{g}$, defined by (15) using this $\lambda$, satisfies Theorem 6.
@@
$g(T)=\frac{1}{R(T)}.$ (53)
The property (6) implies that
$R(KU)\leq K^{\frac{1}{\gamma}}R(U)\,\,\,\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,U>0\,\,\text{and any}\,\,K\,\,\text{large enough.}$ (54)
Applying (54), we see that for $U$ large enough one has
$H\left(\frac{\lambda(U)}{f(U)}\right)=U=R\left(\frac{1}{f(U)}\right)=R\left(\frac{1}{\lambda(U)}\frac{\lambda(U)}{f(U)}\right)\leq\frac{1}{\lambda(U)^{\frac{1}{\gamma}}}R\left(\frac{\lambda(U)}{f(U)}\right).$ (55)
Let us notice that the conditions (6) and (10) on $f$ and (8) on $\lambda$ imply that
$\lambda\left(\frac{\lambda(T)}{f(T)}\right)\asymp\lambda(T),$ (56)
where the implied constant only depends on $\alpha$, $\beta$ and $\gamma$. Combining this with (55), we see that
@@
$H(T)\ll{\lambda\left(T\right)^{-\frac{1}{\gamma}}}R\left(T\right)\,\,\,\,\,\,\,\,\,\,\,\,\text{for all}\,\,T\,\,\text{large enough.}$ (57)
Finally, note that in a similar fashion to (56) one can check that
$\lambda\left(H(T)\right)\asymp\lambda(T),$ (58)
@@
$\text{the set}\,\,\widehat{\bf W}_{n,m}^{\Theta}\left[\frac{1}{4n(2m)^{n}C}u\right]\cap\mathcal{A}\subseteq\widehat{\bf W}^{\Theta}_{n,m}\left[\frac{1}{4n(2m)^{n}}\tilde{g}\right]\cap\mathcal{A}\,\,\text{has measure zero in}\,\,\mathcal{A}.$ (59)
It remains to notice that the function $\lambda$ defined using the function ${4n(2m)^{n}C}\cdot u(T)$ instead of $u(T)$ satisfies conditions (8) and (16), thus (59) holds for $u$ replaced with ${4n(2m)^{n}C}\cdot u(T)$. This completes the proof.
@@ 9 On Khintchine’s result: Theorem D and Theorem 8
Theorem B follows from the following statement, by taking $h(T)=\phi^{n}\left(T^{\frac{1}{m}}\right)$. Statements 1,2 obviously follow form the corresponding statements of the theorem below.
@@
Let $\phi:\,\,[1;+\infty)\rightarrow\mathbb{R}_{+}$ be a decreasing (not necessarily strictly) function.
1. Let $\Theta$ be a real $m\times n$ matrix. For any $\boldsymbol{\eta}\in\mathbb{R}^{n}$, such that the pair $(\Theta,\boldsymbol{\eta})$ is not trivially singular, one has
$\liminf_{t\to\infty}\frac{\psi_{\Theta}(t)}{\phi(t)}\leqslant 2\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(2t)}.$ (60)
2. Let $\Theta$ be a real $m\times n$ matrix. Then, there exists such a $\boldsymbol{\eta}\in\mathbb{R}^{n}$ that
$\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(t)}\leqslant 2B\cdot\liminf_{t\to\infty}\frac{\psi_{\Theta}(t)}{\phi(2A\cdot t)},$ (61)
where $A$ and $B$ are a defined in (17) in Proposition 1.
If in addition if $\Theta$ is not trivially singular, then the set of such $\boldsymbol{\eta}$ is uncountable.
@@
First of all we prove statement 1. of Theorem 1. Let $K=\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(2t)}$. Then for any $\varepsilon>0$ one has
@@
Let ${\bf x}_{\nu}={\bf q}_{n+1}-{\bf q}_{\nu}\in\mathbb{Z}^{m}$. Then $|{\bf x}_{\nu}|\leqslant 2Q_{\nu+1}.$ From the triangle inequality and (62) it follows that the inequality
@@
and (60) is proven.
Now we prove the second statement of Theorem 1.
First, suppose $\Theta$ is trivially singular. Then, one can take $\boldsymbol{\eta}=\Theta{\bf p}$ for some ${\bf p}\in\mathbb{Z}^{m}$; in this case both left and right hand sides of (61) are equal to zero, and the desired statement holds.
@@
We define $\nu_{k}\in\mathbb{Z}_{+},\,k=1,2,3,$... by $P_{\nu_{k}}\leq t_{k}<P_{\nu_{k}+1}$. Then $\psi_{\Theta}(t_{k})=\psi_{\Theta}(P_{\nu_{k}})$ and $\phi(2A\cdot t_{k})\leq\phi(2A\cdot P_{\nu_{k}})$, which implies
@@
$|{\bf y}_{l}|\leqslant A\cdot P_{\nu_{l}}+A\cdot\frac{P_{\nu_{l}}}{2}+\ldots\leqslant 2A\cdot P_{\nu_{l}},$ (63)
@@
Let us take $t\geq|{\bf y}_{l(\varepsilon)}|$. As $|{\bf y}_{l}|\rightarrow\infty$, there exists such an $l$ that $|{\bf y}_{l}|\leqslant t<|{\bf y}_{l+1}|$. From the monotonicity of the function $\psi_{\Theta,\boldsymbol{\eta}}(t)$ and (64) it follows that
@@
  2. B.
    If vectors (5) are linearly independent over $\mathbb{Q}$, then all the terms $\Theta{\bf p}_{\nu_{k}}-{\bf a}_{\nu_{k}}$ are nonzero,

### v3

@@ 7.2 Proof of Theorem 1
Let us consider a decreasing function $\psi$ and function $\rho$ defined by (7.1); notice that such $\psi$ and $\rho$ are related via (7.4). Let
@@
$\sigma(t)=\frac{\varepsilon}{n\rho\left(\frac{m}{\varepsilon}\cdot t\right)}.$ (7.19)
@@
Consider the infinite increasing sequence of integer vectors ${\bf y}\in\mathbb{Z}^{n}$ for which (7.18) holds. We may assume that this sequence is a subsequence of the best approximation vectors for matrix $\Theta^{\top}$. To avoid cumbersome notation, in this proof we do not denote this subsequence as ${\bf y}_{\nu_{k}},k=1,2,3,$..., but simply as ${\bf y}_{\nu},\nu=1,2,3,$... and do not use double indices. Let $Y_{\nu}=|{\bf y}_{\nu}|$. Let
$t_{\nu}=\frac{\varepsilon}{m\psi(Y_{\nu})}.$ (7.20)
It is clear that $t_{\nu}\to\infty$ when $\nu\to\infty$. Therefore, for $\boldsymbol{\eta}\in\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\mathcal{A}$ and for large enough $\nu$ there exists ${\bf q}\in\mathbb{Z}^{m}$ such that $|{\bf q}|\leq t_{\nu}$ and $\Theta{\bf q}-\boldsymbol{\eta}||\leq\sigma(t_{\nu})$.
Now by (7.9), one has
@@
Note that by (7.20) we have
@@
Now the last two equalities by (7.21) imply
@@
From (7.22) we see that
$\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\mathcal{A}\cap\left(B_{M}(0)+{\bf a}\right)\subseteq\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}\Omega_{\nu}$ (7.23)
(recall that now $B_{M}(0)+{\bf a}$ is a $({\rm dim}\,\mathcal{A})$-dimensional ball of radius $M$ in sup-norm in $\mathcal{A}$ centered at ${\bf a}\in\mathcal{A}$), so it is enough to prove that for any fixed $M$ and $\bf a$} the latter set has measure zero under given restrictions on $\varepsilon$. Further, by (7.23) we have
$\left(B_{M}(0)+{\bf a}\right)\setminus\widehat{\bf D}_{n,m}^{\Theta}[\tilde{g}]\supset\left(B_{M}(0)+{\bf a}\right)\setminus\left(\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}\Omega_{\nu}\right)=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{\nu=N}^{\infty}\Omega_{\nu}^{c},\,\,\,\text{where}\,\,\,\Omega_{\nu}^{c}=\left(B_{M}(0)+{\bf a}\right)\setminus\Omega_{\nu}.$
Now, we should note that by the assumption of Theorem 1, $\mathcal{A}$ is non-$\Xi$-exceptional for some $\Xi<1$. This means that (6.1) is not satisfied an so $|{\rm pr}_{\mathcal{L}}\,{\bf y}_{\nu}|\rightarrow\infty$. By Lemma 4 we see that
$m({\Omega}_{\nu})\leq 4\varepsilon(1+\xi(M))\cdot m(B_{M}(0)),$ (7.24)
where $\lim\limits_{M\rightarrow\infty}\xi(M)=0$. As $0<\varepsilon<\frac{1}{4}$, by (7.24) we can take $M$ large enough to ensure the inequality $4\varepsilon(1+\xi(M))<1$. Now
$m({\Omega}_{\nu}^{c})\geq\left(1-4\varepsilon(1+\xi(M))\right)m(B_{M}(0))>Cm(B_{M}(0))$
@@
$\sum\limits_{k=1}^{\infty}m({\Omega}_{\nu}^{c})=\infty.$ (7.25)
As we have noticed in the beginning of our proof, we may take a subsequence of the sequence of all the best approximations. Now it is necessary to take even more sparse subsequence to satisfy the condition
@@
Under this condition we can apply Lemma 5. Together with Lemma J and (7.25) this gives
@@
$m\left(\widehat{\bf D}_{n,m}^{\Theta}\left[\tilde{g}\right]\cap\left(B_{M}(0)+{\bf a}\right)\right)\leq m\left(\bigcup\limits_{N=1}^{\infty}\bigcap\limits_{\nu=N}^{\infty}{\Omega}_{\nu}\right)=m(B_{M}(0))-m(B_{M}(0))=0,$
@@
$\{r:\,\,1\leq r\leq W_{k}^{\prime},\,\,I_{s}({\bf q}^{l}(s))\cap I_{k}({\bf q}^{r}(k))\neq\emptyset\}\leq\Big\lceil\frac{2\kappa\psi_{s}\cdot M_{k}}{2\cdot M_{s}\cdot 2\kappa}+\frac{1}{2}\Big\rceil\leq(1+\varepsilon)W_{k}^{\prime}m\left(I_{s}({\bf q}^{l}(s))\right)$ (7.27)
@@
which implies that $\sum\limits_{k=1}^{\infty}m(E_{k})=\infty$. We also know by (7.27) that for any $s<k$,
@@
The next statement directly implies Theorem 3 via the identification (7.1).
@@
$\text{the inequality}\,\,\,\psi_{\Theta^{\top}}(t)>\psi(t)\,\,\,\text{holds for an unbounded set of $t\in\mathbb{R}_{+}$.}$ (7.28)
@@
First of all, let us notice that the condition (7.28) implies that $\Theta^{\top}$ is not trivially singular (otherwise, $\psi_{\Theta^{\top}}(t)$ would become zero for all large enough $t$), so the sequence of the best approximation vectors for $\Theta^{\top}$ is well defined. Take any $\{\psi_{k}\}$ as in Lemma 9. Let $\{\nu_{k}\}$ be a sequence of natural numbers, such that $\psi_{\Theta^{\top}}(|{\bf y}_{\nu_{k}}|)\geq\psi(Y_{\nu_{k}+1})$, and sparse enough to guarantee $\psi_{k}\frac{Y_{\nu_{k}+1}}{Y_{\nu_{k}}}\rightarrow\infty$.
@@
8 Proof of Theorem 7 and Theorem 4 (c)
We want to deduce Theorem 7 and Theorem 4 (c) from Theorem 6. We will only show the deduction of Theorem 7; the proof of Theorem 4 (c) is purely technical and follows the same lines.
First, note that $\lambda(T):=\left(\frac{u(T)}{g(T)}\right)^{\frac{m\gamma}{m+\gamma n}}$. This implies that condition (3.1) holds.
We recall that $\alpha,\beta$ and $\gamma$ are defined in (2.2), (2.4), (2.5) and (2.6). Properties (2.4) and (2.5) imply that condition (3.2) holds: suppose $\delta>1$ and $T\in\mathbb{R}$, then
@@ 8 Proof of Theorem 7 and Theorem 4 (c)
Thus, the function $\tilde{g}$, defined by (3.3) using this $\lambda$, satisfies Theorem 6.
@@
$g(T)=\frac{1}{R(T)}.$ (8.1)
Property (2.2) implies that
$R(KU)\leq K^{\frac{1}{\gamma}}R(U)\,\,\,\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,U>0\,\,\text{and any}\,\,K\,\,\text{large enough.}$ (8.2)
Applying (8.2), we see that for $U$ large enough one has
$H\left(\frac{\lambda(U)}{f(U)}\right)=U=R\left(\frac{1}{f(U)}\right)=R\left(\frac{1}{\lambda(U)}\frac{\lambda(U)}{f(U)}\right)\leq\frac{1}{\lambda(U)^{\frac{1}{\gamma}}}R\left(\frac{\lambda(U)}{f(U)}\right).$ (8.3)
Let us notice that conditions (2.2) and (2.6) on $f$ and (2.4) on $\lambda$ imply that
$\lambda\left(\frac{\lambda(T)}{f(T)}\right)\asymp\lambda(T),$ (8.4)
where the implied constant only depends on $\alpha$, $\beta$ and $\gamma$. Combining this with (8.3), we see that
@@
$H(T)\ll{\lambda\left(T\right)^{-\frac{1}{\gamma}}}R\left(T\right)\,\,\,\,\,\,\,\,\,\,\,\,\text{for all}\,\,T\,\,\text{large enough.}$ (8.5)
Finally, note that in a similar fashion to (8.4) one can check that
$\lambda\left(H(T)\right)\asymp\lambda(T),$ (8.6)
@@
$\text{the set}\,\,\widehat{\bf W}_{n,m}^{\Theta}\left[\frac{1}{4n(2m)^{n}C}u\right]\cap\mathcal{A}\subseteq\widehat{\bf W}^{\Theta}_{n,m}\left[\frac{1}{4n(2m)^{n}}\tilde{g}\right]\cap\mathcal{A}\,\,\text{has measure zero in}\,\,\mathcal{A}.$ (8.7)
It remains to notice that the function $\lambda$ defined using the function ${4n(2m)^{n}C}\cdot u(T)$ instead of $u(T)$ satisfies conditions (2.4) and (3.4), thus (8.7) holds for $u$ replaced with ${4n(2m)^{n}C}\cdot u(T)$. This completes the proof.
@@ 9 On Khintchine’s result: Theorem D and Theorem 8
Theorem B follows from the following statement, by taking $h(T)=\phi^{n}\left(T^{\frac{1}{m}}\right)$. Statements (a) and (b) of Theorem 8 follow form the corresponding statements of the theorem below.
@@
Let $\phi:\,\,[1;+\infty)\rightarrow\mathbb{R}_{+}$ be a non-increasing function.
  1. (a)
    Let $\Theta$ be a real $m\times n$ matrix. For any $\boldsymbol{\eta}\in\mathbb{R}^{n}$ such that the pair $(\Theta,\boldsymbol{\eta})$ is not trivially singular, one has
    $\liminf_{t\to\infty}\frac{\psi_{\Theta}(t)}{\phi(t)}\leqslant 2\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(2t)}.$ (9.1)
  2. (b)
    Let $\Theta$ be a real $m\times n$ matrix. Then, there exists $\boldsymbol{\eta}\in\mathbb{R}^{n}$ such that
    $\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(t)}\leqslant 2B\cdot\liminf_{t\to\infty}\frac{\psi_{\Theta}(t)}{\phi(2A\cdot t)},$ (9.2)
    where $A$ and $B$ are defined in (3.5) in Proposition 1.
    If in addition $\Theta$ is not trivially singular, then the set of such $\boldsymbol{\eta}$ is uncountable.
@@
First we prove statement (a) of Theorem 1. Let $K=\limsup_{t\to\infty}\frac{\psi_{\Theta,\boldsymbol{\eta}}(t)}{\phi(2t)}$. Then for any $\varepsilon>0$ one has
@@
Let ${\bf x}_{\nu}={\bf q}_{n+1}-{\bf q}_{\nu}\in\mathbb{Z}^{m}$. Then $|{\bf x}_{\nu}|\leqslant 2Q_{\nu+1}.$ From the triangle inequality and (9.3) it follows that the inequality
@@
and (9.1) is proven.
Now we prove (b).
First, suppose $\Theta$ is trivially singular. Then, one can take $\boldsymbol{\eta}=\Theta{\bf p}$ for some ${\bf p}\in\mathbb{Z}^{m}$; in this case both left and right hand sides of (9.2) are equal to zero, and the desired statement holds.
@@
We define $\nu_{k}\in\mathbb{Z}_{+},\,k\in\mathbb{N},$ by $P_{\nu_{k}}\leq t_{k}<P_{\nu_{k}+1}$. Then $\psi_{\Theta}(t_{k})=\psi_{\Theta}(P_{\nu_{k}})$ and $\phi(2A\cdot t_{k})\leq\phi(2A\cdot P_{\nu_{k}})$, which implies
@@
$|{\bf y}_{l}|\leqslant A\cdot P_{\nu_{l}}+A\cdot\frac{P_{\nu_{l}}}{2}+\ldots\leqslant 2A\cdot P_{\nu_{l}},$ (9.4)
@@
Let us take $t\geq|{\bf y}_{l(\varepsilon)}|$. As $|{\bf y}_{l}|\rightarrow\infty$, there exists $l$ such that $|{\bf y}_{l}|\leqslant t<|{\bf y}_{l+1}|$. From the monotonicity of the function $\psi_{\Theta,\boldsymbol{\eta}}(t)$ and (9.5) it follows that
@@
  2. (ii)
    If vectors (2.1) are linearly independent over $\mathbb{Q}$, then all the terms $\Theta{\bf p}_{\nu_{k}}-{\bf a}_{\nu_{k}}$ are nonzero,